OpenStudy (anonymous):

can 110/19 be reduced

OpenStudy (anonymous):

No because 19 is not a prime factor of 110

OpenStudy (anonymous):

on one of my answers for the age thing i got r=110/19 but i need a whole number i think i did this wrong too

OpenStudy (anonymous):

my equation that i got was (r-5)-5=20(r-5)

OpenStudy (anonymous):

and when i tryed to solve i got 110/19

OpenStudy (anonymous):

are you sure it shouldnt be \[-5(r-5) = 20 (r-5)\]

OpenStudy (anonymous):

well the problem says: Gloria is 20 years older than reggie. five years ago, she was five times as old as he was. how old is each now

OpenStudy (anonymous):

The equations you get are g = r+20 g-5 = 5(r-5) =5r-25 Then need to solve these simultaneously ie taking the 2nd - 1st you get -5=4r-45 so 4r=40 r=10 and g=30

OpenStudy (anonymous):

oh i know what i did again shouldnt my first equation be G=R+20

OpenStudy (anonymous):

oh yeah i keep forgetting that part

OpenStudy (anonymous):

20-5 would be 15 right?

OpenStudy (anonymous):

or is it negative